# Network-aided Efficient LLM Services With Denoising-inspired Prompt Compression

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in various tasks, leading to their increasing adoption in diverse services delivered through wireless networks. There is a growing trend toward longer prompts to better leverage LLMs’ capabilities and address difficult tasks. However, longer prompts not only increase data transmission costs but also require more computing resources and processing time, which impacts overall system efficiency and user experience. To address this challenge, we propose Joint Power and Prompt Optimization (JPPO), a framework that combines Small Language Model (SLM)-based prompt compression with wireless power allocation optimization. By deploying SLM at edge devices for prompt compression and employing Deep Reinforcement Learning (DRL) for joint optimization of compression ratio and transmission power, JPPO effectively balances service quality with resource efficiency. Furthermore, inspired by denoising diffusion models, we design a denoising-inspired prompt compression approach that iteratively compresses prompts by gradually removing non-critical information, further enhancing the framework’s performance. Experimental results with long prompt tokens demonstrate that our framework achieves high service fidelity while optimizing power usage in wireless LLM services, significantly reducing the total service response time. With our DRL-based JPPO, the framework maintains fidelity comparable to the no-compression baseline while still achieving a $17\%$ service time reduction through adaptive compression. When prioritizing compression, our framework achieves a compression ratio of up to $16{\rm x}$ while maintaining acceptable fidelity (within a $30\%$ reduction). Compared to no compression, baseline single-round compression with a $16{\rm x}$ compression ratio reduces the system total response time by approximately $42.3\%$, while the denoising-inspired method achieves a $46.5\%$ service time saving.
@@
Large language models, prompt engineering, power allocation, joint optimization
@@ 1 Introduction
Artificial intelligence (AI)’s rapid advancement has become a promising tool for enhancing 6G communication systems [1, 2], providing users with more diverse services. Large language models (LLMs) can offer capabilities to handle various tasks for the mobile users in the network and provide significant advantages in context-aware data processing, real-time inference, and efficient spectrum utilization, making them well-suited for next-generation networks like 6G [3]. Their ability to compress, interpret, and generate relevant information with minimal latency significantly improves the performance of wireless communication systems, especially in resource-constrained environments. Additionally, LLMs have revolutionized natural language processing, demonstrating unprecedented capabilities across various tasks [4]. As these models increasingly drive intelligent IoT devices and edge computing applications, their deployment over wireless networks to provide services to end-users has become a critical scenario [5]. For instance, users can interact with LLM cloud assistants through their mobile devices, sending text queries over wireless networks to remote servers that host the models, which then process and return appropriate responses. However, this integration faces significant challenges from the computational demands of LLMs and the inherent constraints of wireless communication systems [6].
In LLM models, prompts act as input queries that direct the model’s response generation, while tokens are the subword units the model processes to generate text [7]. These tokens are processed through a transformer architecture that consists of multiple layers of self-attention and feed-forward networks [8]. The model first tokenizes the input prompt into a sequence of tokens, which are then embedded and processed through the transformer layers to understand the context and generate appropriate responses. In wireless communication, efficient prompt compression can minimize bandwidth usage while preserving response fidelity. A key issue is the growing length of prompts used to elicit advanced reasoning from LLMs [9], particularly with the advent of techniques like In-Context Learning (ICL) [10] and Chain-of-Thought (CoT) prompting [11]. For example, users may need to upload lengthy documents along with their queries to enable LLMs to provide accurate answers, such as when asking the model to analyze a research paper or respond to a legal document containing thousands of words. While these methods substantially enhance LLM performance, they often necessitate prompts that can extend to tens of thousands of tokens [12]. This trend creates a fundamental trade-off: On the one hand, longer, more sophisticated prompts unlock the full potential of LLMs; on the other, they impose severe burdens on communication bandwidth and computing resources in wireless settings. Moreover, the inference latency significantly increases as longer prompts require more computational resources in the transformer architecture, where each token must attend to all previous tokens, resulting in quadratic complexity growth. This extended processing time delays the generation of the first output token, directly impacting the Quality of Service (QoS) by increasing the initial response time [13]. Consequently, the efficient deployment of LLMs in wireless networks demands innovative solutions that can balance prompt inputs and resource efficiency requirements [14].
@@
Prior research has tried to address these challenges from different angles. The authors in [15] introduced LLMLingua, a coarse-to-fine prompt compression method that demonstrates significant potential for compressing LLM prompts while preserving their semantic integrity. LLMLingua utilizes a small language model for compression, which can be aligned with the target LLM through instruction tuning. While this approach shows promise for reducing communication overhead, it does not fully account for the challenges posed by wireless transmission. The work in [16] proposes LLM-Slice, a system that creates dedicated network slices for LLM services to improve wireless resource management and reduce response delays. Although LLM-Slice advances wireless network architecture for LLM services, it focuses solely on communication resource allocation without considering prompt optimization for individual user requirements. Addressing these issues requires solving two key questions:
@@
  • Q2) How to jointly design wireless resource allocation, particularly transmission power, to meet the latency and power consumption constraints of network-aided LLM services while maintaining high service quality?
@@
For Q1, natural language processing methods offer potential solutions for prompt compression. However, traditional NLP compression techniques, such as text summarization or keyword extraction, often fail to capture the complex reasoning patterns and task-specific requirements embedded in LLM prompts, leading to degraded inference performance. Alternative AI-based approaches, such as large autoencoder models or specialized compression networks, require substantial computational resources and introduce additional inference latency at user terminals, making them impractical for resource-constrained wireless scenarios. Small Language Models (SLMs), which can be easily deployed at user terminals, offer a promising solution through their semantic understanding capability to compress LLM prompts while preserving task-critical information. SLM has been adapted as an effective way to employ transformers for edge applications in [17].
Furthermore, inspired by diffusion denoising models [18], we propose an iterative prompt compression algorithm as a complement to direct SLM-based compression. This approach views the refined, compressed prompt as a “noise-free” state and the original lengthy text as a “noisy” state. Similar to how diffusion models progressively denoise images [18], our algorithm iteratively refines the text through multiple compression stages. This gradual approach addresses a key challenge: while directly compressing a long text might risk losing critical information, performing the compression in stages allows for more controlled information preservation. For instance, achieving a 16x compression through four iterations of 2x compression enables the model to process smaller, more manageable chunks of information at each step, reducing the risk of losing essential content. Despite the iterative nature, the computational overhead remains minimal due to SLM’s efficiency—our experiments show that each compression round only adds about 2% to the total LLM inference time.
@@
Building upon this insight and addressing the Q2, we propose Joint Power and Prompt Optimization (JPPO), a framework that combines SLM-based prompt compression with wireless power allocation optimization, as illustrated in Fig. 1. JPPO captures the trade-off between compression ratio and wireless resource consumption, adapting to both channel conditions and prompt content. This integrated approach enables efficient wireless LLM services while maintaining response quality through intelligent prompt optimization. The contributions of this paper are summarized as
@@
  • We employ a small, aligned language model as a prompt compressor. This approach preserves useful information while significantly reducing the size of the prompt. The SLM efficiently captures essential meaning without requiring training at the transmitter. Additionally, we design a denoising-inspired prompt compression mechanism that performs compression through iterative refinement steps, further improving the system’s overall performance with marginal additional computational overhead.
  • We formulate an optimization problem for our JPPO framework. This problem balances compression quality, wireless transmission performance, and LLM inference efficiency. The objective function is the QoS, considering energy limitation and end-to-end latency, with compression ratio and transmission power as key decision variables.
  • To solve the complex JPPO optimization problem, we employ a Deep Reinforcement Learning (DRL) approach. The DRL agent learns to make optimal decisions on compression ratio and transmit power, effectively reducing communication overhead and accelerating LLM inference services in energy-constrained and variable wireless channel environments.
@@
The remainder of this paper is organized as follows: Section 2 discusses related work in LLM deployment, prompt optimization, and wireless resource allocation. Section 3 presents our system model for wireless network-aided LLM services and introduces the novel denoising-inspired prompt compression method. Section 4 details the JPPO framework, including its mathematical foundations and implementation approach. Section 5 presents comprehensive numerical results and performance analysis. Finally, Section 6 concludes the paper with a summary of our key findings.
@@ 2 Related Work
In this section, we discuss several related works, including LLMs, prompt compression, and wireless network management.
2.1 Large Language Models
The increasing adoption of wireless network-aided LLM-based services has enabled a range of applications, including real-time decision-making and natural language processing services [19]. In [20], the authors proposed a comprehensive vision for designing universal foundation models tailored to the unique needs of next-generation wireless systems and promoted the design of Large Multi-modal Models (LMMs). Additionally, the growth of LLMs and their integration into wireless networks has significantly increased the demand for computational resources and communication bandwidth. Longer prompts exacerbate these challenges, especially in resource-constrained wireless environments, highlighting scalability issues. Efficient resource management is crucial for sustaining AI services in such limited network and computational conditions. In [21], the authors investigated an edge intelligence optimization problem tailored for LLM inference. They formulated an inference model for transformer decoder-based LLMs to maximize inference throughput through batch scheduling and joint allocation of communication and computation resources, while also considering edge resource constraints and varying user requirements for latency and accuracy. The authors in [22] introduced a novel LLM edge inference framework, incorporating batching and model quantization to ensure high-throughput inference on resource-limited edge devices. Then, they formulated an edge inference optimization problem based on the architecture of transformer decoder-based LLMs and solved the problem using the OT-GAH (Optimal Tree-search with Generalized Assignment Heuristics) algorithm. However, existing research on LLM overlooks the high latency caused by the increasing length of prompts in communication systems.
@@ 2.2 Prompt Compression
Prompt compression is a critical technique to mitigate the computational and communication overhead caused by the growing length of input prompts in LLM-based services. Current prompt compression methods face significant challenges due to the limitations of fine-tuning and inefficient compression processes [23]. These challenges include information loss, diminished model capability, and only marginal efficiency gains. In particular, conventional interference management strategies struggle when interference power approaches the level of signal power in wireless communication, leading to a degradation in overall system performance [24]. The authors of [25] proposed Variable Rate Cross-Modal Compression (VR-CMC) and introduced variable rate prompts to represent the data with different grains. In [26], the authors proposed a discrete prompt compression method using Prompt Compression with Reinforcement Learning (PCRL). And it used a computationally efficient policy network to edit prompts. The authors in [23] proposed an innovative framework that strategically tailors LLMs for streamlined context processing by harnessing the synergies among natural language summarization, soft prompt compression, and augmented utility preservation mechanisms. This framework reduces computational overhead and enhances the efficacy of LLMs across various benchmarks. However, prompt compression, combined with wireless power allocation optimization, has not been considered in prior studies. Prompt compression can reduce the data transmission load, while optimizing wireless power allocation ensures that the system operates within the power availability constraints. Jointly optimizing these two factors has the potential to significantly improve the performance and efficiency of wireless systems.
@@ 2.3 Wireless Network Management
Studies in wireless network management have made a significant impact, particularly in power allocation and joint optimization strategies. For power allocation, efficient solutions are crucial for maintaining network reliability under resource constraints. For instance, in [27], researchers developed an online Delayed-Interaction Collaborative-Learning Independent-Decision Multi-Agent DRL (DICLID-MADRL) algorithm that enables access points to independently optimize user selection and power configuration using only local information, thereby enhancing global energy efficiency. Recent studies have also demonstrated the effectiveness of joint optimization approaches in addressing complex wireless communication challenges. In [28], researchers simultaneously optimized multiple parameters, including receive beam-forming, power allocation, High-Altitude Platform Stations (HAPS) positioning, and computation resource distribution using block coordinate descent and successive convex approximation methods. Similarly, in [29], a Graph Neural Network (GNN)-based framework was proposed to jointly solve power control and spectrum allocation problems in a multi-interference environment with shared channels. However, despite these advancements in wireless network management and joint optimization techniques, a significant research gap remains in the context of network-aided LLM services.
In summary, despite the advancements in research discussed above, no existing studies have addressed power allocation strategies in scenarios involving LLM service delivery, particularly in terms of jointly optimizing prompt compression and transmission power for efficient network services. Compared to existing studies, we consider the novel JPPO framework, which combines SLM-based prompt compression with wireless power allocation optimization.

### v4

# JPPO++: Joint Power and Denoising-inspired Prompt Optimization for Mobile LLM Services

@@ Abstract
Large Language Models (LLMs) are increasingly integrated into mobile services over wireless networks to support complex user requests. This trend has led to longer prompts, which improve LLMs’ performance but increase data transmission costs and require more processing time, thereby reducing overall system efficiency and negatively impacting user experience. To address these challenges, we propose Joint Prompt and Power Optimization (JPPO), a framework that jointly optimizes prompt compression and wireless transmission power for mobile LLM services. JPPO leverages a Small Language Model (SLM) deployed at edge devices to perform lightweight prompt compression, reducing communication load before transmission to the cloud-based LLM. A Deep Reinforcement Learning (DRL) agent dynamically adjusts both the compression ratio and transmission power based on network conditions and service constraints, aiming to minimize service time while preserving response fidelity. We further extend the framework to JPPO++, which introduces a denoising-inspired compression scheme. This design performs iterative prompt refinement by progressively removing less informative tokens, allowing for more aggressive yet controlled compression. Experimental results show that JPPO++ reduces service time by $17\%$ compared to the no-compression baseline while maintaining output quality. Under compression-prioritized settings, a reduction of up to 16×16$\times$16 × in prompt length can be achieved with an acceptable loss in accuracy. Specifically, JPPO with a 16×16$\times$16 × ratio reduces total service time by approximately $42.3\%$, and JPPO++ further improves this reduction to $46.5\%$.
@@
Large language models, small language models, prompt engineering, power allocation, joint optimization
@@ 1 Introduction
The rapid advancement of Artificial Intelligence (AI) has positioned it as a key enabler for Sixth Generation (6G) communication systems [1], supporting a broad range of intelligent services. Among AI technologies, Large Language Models (LLMs) have become central to applications such as question answering, code generation, and real-time dialogue [2, 3]. Their ability to interpret and generate context-aware information has significantly advanced natural language processing, enabling a variety of services that are accessed through both cloud and edge environments [4, 5]. As users increasingly access these services via mobile devices, wireless networks have become a critical infrastructure for delivering LLM-powered intelligence [6].
Unlike traditional network-aided services that primarily involve structured data transmission, LLM-based applications impose significantly higher communication demands due to the complexity and variability of natural language prompts and generated responses. Moreover, LLM services often require real-time interaction and are sensitive to prompt transmission delays, presenting new challenges for wireless systems, especially under constraints such as limited transmit power, fluctuating channel conditions, and restricted edge computing resources. These challenges are further amplified in mobile edge scenarios, where users interact with cloud-hosted LLMs via edge devices by sending prompts and receiving responses over wireless links [7]. This mode of service places substantial pressure on the wireless infrastructure, exposing a mismatch between the LLM service requirements and available communications and computing resources. Addressing this mismatch requires new system-level optimization frameworks that adaptively manage transmission and processing strategies to maintain service quality under dynamic network conditions [8, 9].
Among the various factors contributing to the system load, the structure and length of the input prompt play a central role. In LLMs, a prompt is first tokenized into a sequence of subword units, known as tokens [10]. These tokens are embedded and passed through a transformer architecture comprising multiple layers of self-attention and feed-forward networks [11], which model contextual dependencies and generate the final output. This computation scales with prompt length, as each token attends to all previous tokens, resulting in quadratic time and memory complexity. Furthermore, recent prompting strategies such as In-Context Learning (ICL) [12] and Chain-of-Thought (CoT) [13] have pushed LLMs toward more advanced reasoning, but often at the cost of significantly longer prompts [14]. In practical scenarios, users may upload entire documents along with queries, such as research papers or legal texts containing thousands of words, to elicit accurate and context-rich responses. While these techniques improve LLM service quality, they commonly result in prompt lengths reaching tens of thousands of tokens [15]. This introduces a fundamental trade-off: longer, more informative prompts enhance output quality but impose heavy burdens on wireless transmission and edge-side inference.
@@
Prior research has approached these challenges from both the model and network perspectives. From the model side, the authors in [16] introduced LLMLingua, a coarse-to-fine prompt compression method that demonstrates significant potential for compressing LLM prompts while preserving their semantic integrity. LLMLingua utilizes a Small Language Model (SLM) for compression, which can be aligned with the target LLM through instruction tuning. While this approach shows promise for reducing communication overhead, it does not fully account for the challenges posed by wireless transmission. From the network side, LLM-Slice [17] introduces a wireless slicing mechanism that reserves dedicated network resources for LLM services, thereby reducing response latency. Yet this method focuses purely on communication resource allocation and does not incorporate prompt-level adaptation tailored to individual service requirements. These limitations highlight the need for a unified solution to address two key questions:
@@
  • Q2) How to jointly allocate wireless resources, particularly transmission power, to meet the latency and energy constraints of network-aided LLM services?
@@
For Q1, natural language processing methods offer potential solutions for prompt compression. However, traditional NLP compression techniques, such as text summarization or keyword extraction, often fail to capture the complex reasoning patterns and task-specific requirements embedded in LLM prompts, leading to degraded inference performance. Alternative AI-based solutions, including large autoencoder models or task-specific compression networks, typically demand substantial computational resources and introduce additional latency at the user side, rendering them unsuitable for wireless scenarios with constrained devices. Motivated by LLMLingua [16], we consider SLMs as a practical solution for prompt compression to achieve resource-efficient deployment. SLMs can be easily integrated into user terminals and offer sufficient semantic understanding to retain task-critical content during compression. They have also been shown to be effective in enabling transformer-based inference in edge applications [18].
Furthermore, inspired by the diffusion models, especially the Denoising Diffusion Probabilistic Models (DDPM) [19], we propose an iterative prompt compression algorithm to complement direct SLM-based compression. In this approach, the original long prompt is treated as a “noisy” input, and the refined compressed version represents the “denoised” state. Analogous to how diffusion models progressively remove noise from images, our method performs multi-stage prompt refinement, gradually filtering out redundant or non-essential content. This design mitigates the risk of semantic loss that can occur when compressing lengthy prompts in a single step. This principle is similar to how humans revise long-form text: reducing a complex document by $90\%$ in one attempt often results in missing key points, while multiple rounds of revision, each removing a manageable portion, allow for better content preservation and clarity. Likewise, a 16×16$\times$16 × compression can be achieved through four iterations of 2×2$\times$2 × compression, enabling the model to operate on smaller, semantically coherent segments at each stage. This controlled process enhances the likelihood of retaining task-critical information while maintaining high compression. Despite its iterative structure, the method introduces minimal computational overhead due to the efficiency of SLMs. Empirically, each compression round adds only about $2\%$ to the total LLM inference time, making it well-suited for latency-sensitive wireless applications.
@@
Building upon this insight and for addressing the Q2, we propose Joint Power and Denoising-inspired Prompt Optimization (JPPO++), a framework that combines SLM-based prompt compression with wireless power allocation optimization, as illustrated in Fig. 1. JPPO++ captures the trade-off between compression ratio and wireless resource consumption using Deep Reinforcement Learning (DRL), adapting to both channel conditions. The contributions of this paper are summarized as
@@
  • We utilize a small, aligned language model as a prompt compressor in JPPO++ for network-aided LLM services. The SLM efficiently captures essential meaning while significantly reducing the length of the prompt without requiring training at the transmitter. Additionally, we design a denoising-inspired prompt compression mechanism that performs compression through iterative refinement steps, further improving the system’s overall performance with marginal additional computational overhead.
  • We formulate a joint optimization problem that captures the trade-offs among wireless transmission efficiency and LLM service quality. The objective is to maximize end-to-end Quality of Service (QoS) under energy and delay constraints, where the compression ratio and transmission power serve as coupled decision variables.
  • To solve this dynamic optimization problem, we employ a Deep Reinforcement Learning (DRL) framework. The DRL agent learns adaptive compression and power strategies based on real-time network conditions and prompt characteristics. This enables JPPO++ to outperform static baselines by reducing communication cost and improving responsiveness in resource-constrained wireless environments.
@@
The remainder of this paper is organized as follows: Section 2 discusses related work in LLM deployment, prompt optimization, and wireless resource allocation. Section 3 presents our system model for wireless network-aided LLM services and introduces the novel denoising-inspired prompt compression method. Section 4 details the JPPO++ framework, including its mathematical foundations and implementation approach. Section 5 presents comprehensive numerical results and performance analysis. Finally, Section 6 concludes the paper with a summary of our key findings. A list of mathematical symbols frequently used in this paper is shown in Table I.
TABLE I: Mathematical Notations
Notation Description
${\textit{{x}}}_{\text{ins}},{\textit{{x}}}_{\text{dems}},{\textit{{x}}}_{\text{que}}$ The instruction component, the demonstrations or examples, and the specific question or task that consisted in an original prompt x
$\hat{\textit{{x}}}$ The compressed prompt
$\mathcal{L}_{x},\mathcal{L}_{\hat{x}}$ The total token length of the original prompt and the token length of the compressed prompt
$\mathbf{f_{1}},\mathbf{f_{2}},\mathbf{f_{3}}$ The three essential aspects of semantic preservation during information transmission to evaluate the quality of prompt compression
$\phi_{1},\phi_{2},\phi_{3}$ The weight factors determining the relative importance of each fidelity component
$\alpha(t)$ The compression ratio at any time $t$
$\psi_{\kappa}$ The target compression ratio
$\sigma(t)$ A monotonically increasing scheduling function
$\beta(i)$ The compression ratio between consecutive steps
$E_{e}$, $E_{t}$ The encoding energy and transmission energy
$R$ The transmission rate
$w$ The bandwidth of the offloading link between the user and DCO
$\gamma$ The Signal-to-Noise Ratio (SNR)
$\omega$ The path-loss exponent
$\lambda^{2}$ The Gaussian noise term in the Additive White Gaussian Noise (AWGN) channel
$T$ The total time consumption
$t_{\text{e}}^{{\rm SLM}}$, $t_{\text{e}}^{{\rm LLM}}$, $t_{\text{t}}$ The encoding delay in both SLM and LLM, and transmission delay
$\varphi$ The Probability Density Function (PDF) expression for wireless channel fading
$P_{\text{T}}$ The transmit power
$\eta$ BEP
$P_{\text{th}}$ The maximum allowable power consumption
$T_{\text{th}}$ The maximum tolerable end-to-end latency
$\mathbf{f}_{\text{th}}$ The minimum required fidelity
@@ 2 Related Work
This section discusses several related works about LLMs, prompt compression, and wireless network management.
2.1 Large Language Model Services
The integration of LLMs into wireless network-aided services has enabled a range of intelligent applications, including real-time decision-making and natural language understanding [20]. However, this trend imposes increasing demands on both computation and communication resources due to the growing size of LLMs and the complexity of their input prompts. To address computational bottlenecks, several recent efforts have focused on optimizing LLM inference at the network edge. For example, the authors in [21] formulate a scheduling problem that jointly allocates computing and transmission resources for transformer-based LLMs, maximizing throughput under edge constraints. Similarly, the authors in [22] propose an edge inference framework that combines model quantization and batching, and develop the Optimal Tree-search with Generalized Assignment Heuristics (OT-GAH) algorithm for efficient scheduling on resource-limited devices. These studies highlight the importance of inference-side optimization, yet they operate under an implicit assumption that the input prompt is already available. In practice, especially in cloud-based LLM deployments accessed via wireless networks, uploading long prompts often accounts for a substantial portion of the total latency. This issue is exacerbated by recent prompting strategies, such as ICL and CoT reasoning, which require longer, more structured inputs to fully exploit model capabilities. As prompt length grows, communication costs become a dominant factor in end-to-end performance, particularly in bandwidth-limited or energy-constrained environments. While the authors in [23] offer a broad vision for aligning foundation models with the needs of wireless systems and discuss emerging Large Multi-modal Models (LMMs), they do not directly address prompt transmission overhead or latency. This gap highlights a critical insight: without coordinated control of both prompt input and transmission behavior, LLM services over wireless networks will remain bottlenecked by front-end communication delays. Thus, prompt compression and adaptive resource control must be treated as essential optimization targets, jointly considered alongside inference design to support scalable, low-latency LLM services in next-generation networks.
@@ 2.2 Prompt Compression
Prompt compression has emerged as a crucial technique for mitigating the computing and communication overhead introduced by increasingly lengthy prompt inputs in LLM services. However, existing compression methods face significant challenges, including limited compression efficiency, reliance on fine-tuning or task-specific heuristics, and the risk of degrading LLM performance due to loss of critical contextual information [24]. These limitations are particularly critical in wireless environments, where bandwidth and energy constraints amplify the impact of prompt length. Moreover, conventional interference management techniques become less effective when signal and interference power levels converge, further degrading system-level performance [25]. Several recent works have attempted to improve prompt processing efficiency. VR-CMC [26] introduces variable-rate prompts for cross-modal compression, allowing data to be represented at multiple levels of granularity. PCRL [27] formulates prompt editing as a discrete decision process and applies reinforcement learning to learn efficient compression policies. Furthermore, the authors in [24] propose a framework that integrates summarization, soft prompt compression, and utility-aware mechanisms to reduce context processing load while maintaining model utility across tasks. While these approaches offer meaningful gains in compression quality and model efficiency, they are primarily designed without considering the dynamics of wireless systems. Crucially, none of these methods consider the joint effects of prompt compression and wireless transmission control. In bandwidth and energy-constrained settings, compression alone reduces data volume but cannot guarantee reliable or timely delivery. Conversely, adaptive power allocation improves link quality but does not reduce the underlying transmission load. These two levers—prompt compression and wireless resource control—are inherently complementary. Their joint optimization is crucial for enabling efficient, low-latency LLM services over wireless networks, particularly under stringent QoS and energy constraints.
@@ 2.3 Wireless Network Management
Extensive research in wireless network management has led to effective solutions for power allocation and joint optimization under resource constraints. For example, the authors in [28] propose the Delayed-Interaction Collaborative-Learning Independent-Decision Multi-Agent DRL (DICLID-MADRL) algorithm, where access points independently optimize user selection and power configuration using only local observations, improving global energy efficiency through decentralized multi-agent reinforcement learning. Beyond power control, recent studies have addressed more complex joint optimization problems. In [29], a block coordinate descent and successive convex approximation framework is applied to jointly optimize receive beamforming, power allocation, HAPS positioning, and computation resource distribution. Similarly, the authors in [30] employ a graph neural network-based method to jointly solve power control and spectrum allocation in multi-user interference scenarios with shared channels. These works demonstrate the potential of learning-based and algorithmic strategies for managing wireless resources in complex environments. However, they do not account for the unique characteristics of network-aided LLM services, where large, dynamic prompt inputs and strict latency constraints fundamentally alter the nature of the optimization problem. In particular, the interplay between input compression and transmission power remains unaddressed. Bridging this gap requires extending joint optimization frameworks to explicitly incorporate the semantic structure and variability of LLM inputs alongside traditional physical-layer resource control.
In summary, while prior studies have made significant progress in both prompt compression and wireless resource management, they treat these aspects in isolation. Our work departs from this separation by introducing JPPO++, which enables efficient and scalable LLM service delivery in resource-constrained wireless networks.